OpenStudy (anonymous):

Determine the area of the enclosed by y=3x2 and y=6x and Determine the area enclosed by the graphs of y=x3 and y=4x for x≥0 what is really the difference between the two questions and how are they solved

OpenStudy (xapproachesinfinity):

this is not double integral no?

OpenStudy (xapproachesinfinity):

what is the interval for the first one

OpenStudy (anonymous):

nno intervals given but with the interception, we can find the intervals

OpenStudy (xapproachesinfinity):

yes correct! what did you find from the graph

OpenStudy (anonymous):

wish is x1=0 to x2=2 for the first one

OpenStudy (xapproachesinfinity):

so \(\large \int_{0}^{2}(3x^2-6x)dx\)

OpenStudy (xapproachesinfinity):

area enclosed by the curves

OpenStudy (anonymous):

yes

myininaya (myininaya):

Is 6x>3x^2 on [0,2]?

OpenStudy (xapproachesinfinity):

actually that depend on which is upper and lower!

OpenStudy (xapproachesinfinity):

so could be 6x-3x^2

myininaya (myininaya):

It is 6x-3x^2 since 6x>3x^2 on [0,2]

OpenStudy (anonymous):

my result gave 4.... but when i tried solving the second question, i got different result

OpenStudy (anonymous):

yes @ myininaya

OpenStudy (xapproachesinfinity):

yeah right it is

OpenStudy (anonymous):

can you help with the second

OpenStudy (xapproachesinfinity):

well it is the same process! what are your boundaries

OpenStudy (anonymous):

no interval

OpenStudy (xapproachesinfinity):

well there must be

OpenStudy (xapproachesinfinity):

[0,?]

OpenStudy (anonymous):

0,2

OpenStudy (anonymous):

am just lost in this..... the answer to the first one is 8 and to the second, it is 4..... please enplane,,,,, what is the difference

OpenStudy (xapproachesinfinity):

so 4x>=x^3 in the interval [0,2]

OpenStudy (xapproachesinfinity):

this is what you did

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

and you got 4

OpenStudy (anonymous):

yes after integrating and substituting x1 and x2

OpenStudy (xapproachesinfinity):

do the regions looks similar?

OpenStudy (anonymous):

that is what i do not get

OpenStudy (xapproachesinfinity):

i haven't done the graphs so don't know

OpenStudy (anonymous):

ok. thanks for your help any ways

OpenStudy (xapproachesinfinity):

you are welcome! i forgot this stuff lol

myininaya (myininaya):

@GIL.ojei Are you saying you are having trouble with the integration?

OpenStudy (anonymous):

well, yes but mostly,how to know the region

myininaya (myininaya):

Well it is obvious to me you can find the intersections pretty easily...You have been able to answer that question each time.

myininaya (myininaya):

If you want to know what function is larger than the other, you could test the interval in between each intersections by pluggin in a number

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

i think he didn't get which function is larger than the other in the interval

myininaya (myininaya):

Say I want to find the area bounded by y=x^2 and y=x. So x^2=x when x=1 or x=0 . We know x>x^2 because 1/2>1/4 (when you plug in 1/2) so we know we are going to have to do int(x-x^2,0..1)

OpenStudy (xapproachesinfinity):

you can see it clearly from the graph

myininaya (myininaya):

You could also do this with a visual as @xaspprachesinfinity suggested However, if you want to skip the visual you can. But a visual does help.

OpenStudy (anonymous):

@xapproachesinfinity ,i do not know how to draw the graph

myininaya (myininaya):

@GIL.ojei Do you understand what I was saying? You know how to find the intersections for y=x^3 and y=4x for x>=0 you seen that x^3=4x when x=-2,0,2 x=-2 does matter since x>=0 so we are only looking as a visual between 0 and 2 for the graphs y=x^3 and y=4x To see which is bigger on that interval, simply plug in a number between 0 and 2 like 1. Let f(x)=x^3 and g(x)=4x f(1)=1 g(1)=4 g>f since 4>1 so 4x>x^3 on the interval [0,2]

OpenStudy (xapproachesinfinity):

you draw both graphs you will see that they intersect somewhere that the same thing as saying 3x^2=6x to find the intersections

myininaya (myininaya):

so you do big-small

myininaya (myininaya):

doesn't matter* (i didn't mean to say does matter)

OpenStudy (xapproachesinfinity):

i find graphs to be much more easier to understand! Ate any rate, the result is the same^_^

myininaya (myininaya):

Graphs are very useful.

OpenStudy (anonymous):

so if x1 > x2 what becomes the region @myininaya

myininaya (myininaya):

So I would learn how to graph these basic algebraic graphs.

myininaya (myininaya):

what is x1 and x2?

OpenStudy (anonymous):

the intervals

myininaya (myininaya):

 If the function f is bigger than the function g on the interval [x_1,x_2], then you setup the integral as follows: \[\int\limits_{x_1}^{x_2}(f(x)-g(x)) dx \]